OpenStudy (anonymous):

Find the area of a regular hexagon with radius 12 in. Round to the nearest whole number.

OpenStudy (blurbendy):

A = \[\frac{ 3\sqrt{3}*r^2 }{ 2 }\]

OpenStudy (blurbendy):

where r is the radius

OpenStudy (anonymous):

Thank you both! Appreciate it =] @blurbendy @ybarrap

OpenStudy (ybarrap):

Regular hexagons have 6 equilateral triangles, each with height (h) equal to it's apothem: h = apothem = sqrt(r^2 - (r/2)^2) A = 6 * (1/2) base * h = 6 * (1/2) * r * r * sqrt(3)/2 = 3*sqrt(3)*r^2/2 where r is radius (12 in).

OpenStudy (blurbendy):

yw :)

OpenStudy (ybarrap):

np